1.The Perimeter of a rectangle is 60cm. Its length is twice its width. Find its length. Thank you so much.
Answer by ewatrrr(24785)

 
Hi,

length is twice its width
Let x and (2x) represent the width and length respectively.
Question states***
P = 2*2x + 2x = 60cm
6x = 60cm
x = 10cm, the width. the length is 20cm
CHECKING our Answer***
2*10cm + 2*20cm = 60cm